# A scanning apparatus to detect the spectral directivity of optically-emissive materials

### v1

@@
A scanning apparatus to detect the spectral directivity of optically-emissive materials
††journal: opticajournal††articletype: Research ArticleAn apparatus that records the optical spectrum of emissive materials as a function of the polar coordinate angles is reported. The ability for the device to characterize the directivity of a light source over the optical spectrum is demonstrated. The angular dependence of an electrically-driven LED with a hemispherical diffuser cap was characterized. Optically pumped materials that emitted both fluorescence and amplified spontaneous emission (ASE) were also characterized.
1 Introduction
The passive linear optical properties of a material can be characterized by detecting changes to an incident probe beam using methods such as reflection, transmission, and absorption spectroscopy.[1] Active techniques involve probing the material upon exciting it with i) an external light source,[2] ii) electrically-driven charge carriers,[3] or iii) chemical reactions.[4] Measuring the directivity over a range of the light spectrum exiting a sample after relaxing from an excited state is the primary focus of this paper.
@@ 1 Introduction
When a high concentration of fluorophores is pumped by a strong light source, a population inversion density can be achieved. Many fluorescent molecules behave like a 4-level model.[9, 10] If the path length of an emitted photon through the gain region of excited molecules is long enough, then amplified spontaneous emission (ASE) can occur.[11] Traditionally, a long strip of material is excited to produce ASE.[12, 13] Aperiodic scattering surfaces inside the gain region can increase the effective path length of emitted photons which can create directed ASE.[14, 15, 16, 17] The narrow spectral profile for this type of enhanced ASE generation is similar to the spectrum produced by multimode laser emission from an “intensity feedback” or “non-resonant feedback” random laser.[18, 19, 20, 21] Natural and fabricated materials can possess a highly periodic internal structure.[22, 23, 24, 25, 26, 27] When the internal structure is highly periodic, laser light can be emitted (e.g., multilayer distributed feedback laser).[28, 29, 30, 31] At high thresholds, even aperiodic structures can produce laser emission, which has been referred to as a “resonant feedback” random laser.[32, 33, 34] The directivity from the complex structures and properties given in the latter examples can be probed by measuring the spectrum of the emitted light over the polar and azimuth angles. This paper presents a spectral scanner, which collects light via a miniature spectrometer while scanning through the polar and azimuth angles to characterize the emissive properties of simple and complex-structured materials and devices.
2 Experimental methods
2.1 Apparatus
@@
Automated recordings of emission spectra scanned over the azimuth and polar angles were performed by an in-house built apparatus. An aluminum sample stage was at the center of the apparatus and temperature controlled by running liquid through a bored cavity. The sample stage was secured by a post that ran through a bore hole in a rotation stage that swept the azimuth angles as shown in Figure 1. A circular plate was attached to the rotation stage. An off-center rotation stage was attached to sweep the polar angles. The circular plate was used to mark rotation angles to initially configure the azimuth rotation stage, but a horizontal jib arm holding the second rotation stage was also found to work in the same capacity. A boom arm was attached to the second rotation stage. A lever arm was attached at the top of the boom arm to secure a sensor and its alignment components.
@@
A tunable double lens tube was used to zoom and focus in on the sample stage. Adjustments to the lens tube’s orientation were made by mounting it on a kinematic mirror mount. The kinematic mirror mount is mounted on a slot and the distance between zoom lens tube and the sample plate can be adjusted, A camera was placed in the image location for alignment purposes. After alignment, the camera was replaced by a m-core multimode optical patch fiber. The fiber was connected to a portable spectrometer mounted to the boom arm. We initially attempted to secure one end of a long fiber to the apparatus and the other end to a spectrometer mounted on the optical table, but fiber deformations from apparatus rotations caused the relative intensity of the signal to change. By mounting the spectrometer to the boom arm with a rigid fiber patch cord, loss through the optical fiber was held constant. A slot was located between the lens tube and the detector position which allowed for 1" filters (chromatic, polarization, etc.).
@@
The motors on the apparatus were controlled by a CNC shield attached to an Arduino Uno loaded with a GRBL library. The Arduino was connected to a computer via USB port. The spectrometer was also connected to the computer. A Python script recorded spectra with the Seabreeze package, and it also controlled the apparatus motors. A third motor was connected to the CNC shield which rotated a halfwave plate as shown in Figure 1. The pump beam’s polarization was rotated and sent through a polarizing beam cube splitter. Pulse energy was adjusted using this halfwave plate and polarizer combination. A relay was also connected by USB port and incorporated into the same Python code that controlled the entire apparatus. The relay controlled a shutter, where the shutter was open for data collection of materials excited by the external pump laser and closed when the apparatus rotated during the scanning procedure. The shutter reduced fluorophore photodegradation during a scan.
2.2 Alignment
@@ 2.3 Experiment
The apparatus measures the directivity of an emissive material/device over the spectral range recorded by the fiber spectrometer. Two types of samples were recorded, 1) an electrically-driven light emitting diode (LED) with a diffuser cap and 2) a fluorescein-doped water droplet. The electrical current for the LED and the pump beam pulse energy for the fluorescein-doped water were held constant for the duration of their scans. All scans were stepped in the azimuth from to . They scans were stepped along the co-latitude from to .
@@
The two different sizes of droplet samples were measured under the same geometric condition. For the first one, a large dye-doped water droplet was placed on a hydrophobic parafilm surface with a pump beam illuminating the sample from above. A second droplet was approximately 4 times smaller than the first droplet and tested by the similar size pump beam. A nm pulsed laser beam with a ns duration was used for this experiment.
3 Results
3.1 Diffuser-capped LED
@@
Our system was aligned through the process described in Section 2.2. In an attempt to confirm the alignment, an LED with a diffuser cap was placed at the center of the sample stage. An ideal spherical diffuser cap would mimic a spherically symmetric light source, which would confirm the alignment; however, the spherical diffuser capping the LED did not have spherical light-emissive symmetry.
@@
At high polar angles, a hot spot was formed along the azimuth as depicted in the intensity heat map shown in Figure 2(a). The nonuniform directivity could have been caused by an alignment issue, so the sample was rotated by and a new scan was performed. As shown in Figure 2(b), the recorded hot spot rotated in the azimuthal direction by the same angle that the sample was rotated. This azimuth rotation of the sample confirmed the hot spot was indeed sourced from the diffuser-capped LED as opposed to an issue with the system alignment.
@@
As expected, there was no distinguishable difference between the peak wavelength and spectral width of the LED’s emitted light. The amplitude of the spectra shown in Figure 2(c-f) vary in amplitude only. There is a significant trend for the intensity’s dependence on the polar angle as shown in Figure 2(c-d). The detected signal is largest at near the top of the LED and significantly weakens as the detector collects light near the horizon. A hot spot is clearly identifiable as a function of the azimuthal angle. The peak amplitude of the spectra, shown in Figure 2(e-f), shifted by the same angle as the rotated sample. The LED has no significant process for emitting directed energy, and the diffuser cap is fabricated from a passive achromatic material. Therefore, as expected for a linear optical source, the spectral features remain constant while only the amplitude of the emitted light depends on the spherical angles. This result implies that no directed energy occurs at the source or during the passive diffusion through the cap, as expected.
3.2 Large water droplet: center pumped
@@
The diffuser-capped, electrically-pumped, LED light source in the previous section had hot spots from aspherical light emission. A wt.% fluorescein/water solution was used in an attempt to create a better approximation to a spherical-emitting light source. The droplet was placed on parafilm, where the solution/air interface above the parafilm surface formed an approximate hemisphere. The beam spot was in diameter, which illuminated the majority of the liquid hemisphere’s surface. The liquid cooled stage was measured to be C while the lab was at a typical laboratory room temperature and humidity. The cooled surface resulted in a near steady state droplet size due to the competition between evaporation and condensation.
@@
There is an increased emission intensity at higher co-latitude (lower latitude) for the droplet. This trend in emission as a function of polar angle is opposite of the diffuser-capped LED, where the spectra shown in Figure 3(a) can be visually compared those given in Figure 2(c).
@@
The noise associated with the spectrum is also greater in the droplet at low emission intensity levels as compared to the diffuser-capped LED. The azimuthal dependence shown in Figure 2(b) also has some shot-to-shot variations throughout the scan. Some small variations in the beam profile could result in an azimuth-dependent emission; however, shot-to-shot noise from the pump beam also contribute to the variations in the azimuthal dependence of the emission intensity. The shot-to-shot noise of an Nd:YAG’s fundamental frequency is reasonable low. The shot-to-shot uncertainty of the pump is increased when a portion of the nm pump beam’s photons are frequency doubled by a beta Barium Borate (BBO) crystal followed by sum-frequency generation of the nm and nm to yield a nm beam. This nonlinear, macroscopic, cascading process results in larger variations between pulse energies for the nm beam entering an optical parametric oscillator (OPO). For the OPO stage, the residual nm and nm light is attenuated by a set of dichroic mirrors. The remaining nm light is injected into a cavity via another dichroic mirror with a long BBO crystal between two mirrors. The amplified output of the OPO for wavelength selection at nm also amplifies the shot-to-shot uncertainty of the pulse energy. Thus, the uncertainty between each shot appears to be relatively large.
@@
Averaging out the uncertainty between recorded shots taken while the azimuth is scanned, a slow change of intensity is observed through the scan as shown in Figure 3(b). The output intensity has some preferred azimuthal emission direction. If the preferred direction were biaxial, then the cause would undoubtedly be from the pump beam’s polarization. The excited state lifetime of fluorescein is relatively large to the pump pulse duration, making a polarization dependent emission unlikely. Indeed, we found that the small, preferred, azimuth emission was unidirectional, which was caused by a uncertainties in the profile and center-alignment of the OPO pump beam. Although the overall signal is relatively small compared to the single-shot noise of the spectrometer, there is a clear direction of preferred emission. The ASE clearly dominates the fluorescence signal. The small fluorescence peak is observed at a nm wavelength while the peak of the red-shifted ASE profile is much greater at nm.
3.3 Small water droplet: off-center pumped

### v2

# Scanning apparatus to detect the spectral directivity of optically-emissive materials

@@
Abstract
An apparatus that records the optical spectrum of emissive materials as a function of the polar coordinate angles is reported. The ability for the device to characterize the directive gain of a light source over the optical spectrum is demonstrated. The angular emission profile of an electrically driven LED with a hemispherical diffuser cap was measured. In addition, the device was used to characterize optically pumped materials exhibiting both fluorescence and amplified spontaneous emission (ASE), demonstrating its versatility for diverse emissive systems.
I Introduction
Apparatus designed to detect the directive gain of an emissive material or device are most commonly applied to antennas. Directive gain measurements of transmitting antennae actuate a rotation of the antennae about the azimuth and polar angles while keeping the receiving antenna static.[1] It is also standard to orient a transponder antennae that can both transmit and receive signals while using a static mirror to bounce the signal from a specific location to determine the directive gain. Anechoic chambers to dampen outside influence are commonplace in antenna frequency ranges.[2] For electrically powered optical components, the same antenna directive gain measurement system can be used; however, in optically pumped systems such as materials doped with fluorophores (fluorescent molecules), the pump orientation with respect to the sample should remain static. Therefore, we have designed a system that maintains the sample at static orientation in the lab reference frame while sweeping the polar and azimuthal angles with the detector. Furthermore, materials and devices that emit at optical frequencies often have broad spectral output with combined processes of fluorescence, amplified spontaneous emission (ASE), and laser emission that can be highly directional. Therefore, a directive gain measurement apparatus that scans over the visible spectrum with a miniature spectrometer was developed. The sample and pump beam remains in a fixed position in the lab frame, and the detector sweeps the azimuth and polar angles measuring angular dependent spectra.
This study investigates the directive gain of light emitted from a material for a given sample orientation. Such materials and techniques probe the “active” emissive properties of the material. Passive linear optical properties of a material are characterized by detecting changes to an incident probe beam using methods such as reflection, transmission, and absorption spectroscopy.[3] Active techniques involve probing the material upon exciting it with i) an external light source,[4] ii) electrically-driven charge carriers,[5] or iii) chemical reactions.[6] Measuring the directivity over a range of the light spectrum exiting a sample after relaxing from an excited state is the primary focus of this paper.
@@ I Introduction
When a high concentration of fluorophores is pumped by a strong light source, a population inversion density can be achieved. Many fluorescent molecules behave like a 4-level model.[11, 12] If the path length of an emitted photon through the gain region of excited molecules is long enough, then amplified spontaneous emission (ASE) can occur,[13] which is light emitted with a high degree of spatial coherence but without temporal coherence. ASE is easily identified in emission spectra by the nonlinear power dependence of the emission intensity and spectral width. Traditionally, a long strip of material is excited to produce ASE.[14, 15, 16, 17] Aperiodic scattering surfaces inside the gain region can also increase the “effective” path length of emitted photons which can create directed ASE.[18, 19, 20, 21] The narrow spectral profile for this type of enhanced ASE generation is similar to the spectrum produced by multimode laser emission from an “intensity feedback” or “non-resonant feedback” random laser.[22, 23, 24, 25] Natural and fabricated materials can possess a highly periodic internal structure.[26, 27, 28, 29, 30, 31] When the internal structure is highly periodic, laser light can be emitted (e.g., multilayer distributed feedback laser).[32, 33, 34, 35] At high thresholds, even aperiodic structures can produce laser emission, which has been referred to as a “resonant feedback” random laser.[36, 37, 38] The directive gain from the complex structures and properties given in the latter examples can be probed by measuring the spectrum of the emitted light over the polar and azimuth angles. This paper presents a spectral scanner, which collects light via a miniature spectrometer while scanning through the polar and azimuth angles to characterize the emissive properties of simple and complex-structured materials and devices.
II Preliminary theory
To understand scan results from the apparatus, it may be beneficial to first discuss the directive gain of a radiative source. The time-averaged power density (the intensity) vector of a monochromatic wave can be determined from the Poynting theorem,
with the asterisk denoting the complex conjugate. We have expressed the electric field and magnetic field in the usual phasor form for a monochromatic wave of angular frequency , where
Here, the field in phasor form must follow as
where is a constant that relates the initial phase of the wave when .
In the far-field, the radiated power is directed radially outward from the source, . Therefore, alignment of the sample to be the same distance away from the detector at all scanned angles is crucial to the accuracy of the scan. The average power radiated is found by integrating the intensity by the area of a sphere centered at the source in the far field. The intensity for some small solid angle patch is maximized for some set of spherical coordinate angles, . The normalized radiation intensity as a function of the spherical polar coordinates is expressed as
It is often useful to find the average value for the normalized radiation intensity over an entire sphere with the source at the center. In this case,
where is the solid angle. The denominator in Equation 5 is . For a monochromatic source such as an antenna, the directive gain is determined by the ratio of the normalized radiation intensity and the average normalized radiation intensity integrated over an entire sphere,
The directivity is defined as the maximum directive gain. A simple way to determine the directivity of a source emitting a monochromatic wave is by first determining the pattern solid angle,
Note that Equation 7 allows the average normalized radiation intensity integrated over the entire sphere to be rewritten, where . From the normalization condition, it also follows that the maximum value of is 1. Therefore, the directivity, which is the maximum value of the directive gain, can be easily expressed in terms of the pattern solid angle,
As written in Equation 8, the directive gain is defined for monochromatic sources. Many visible light sources have a broad emission bandwidth. Although the data is typically graphed with the vertical axis labeled as “intensity,” spectrometers typically measure the intensity spectral density, , which has dimensions of intensity per unit frequency or per unit wavelength. Integrating the spectral curve yields the total intensity of light, i.e., . For a spectrometer measuring the intensity spectral density over the spherical angles, the stored data is in the form of a hypercube. This requirement is akin to storing hyperspectral image data that stores the Cartesian () image positions as well as the multiple wavelength channels.
We introduce the wavelength-dependent, spectral directive gain density, . Integrating the wavelength-dependent, spectral directive gain density over all wavelengths yields the directive gain for the total power emitted over the entire spectrum,
The normalization condition requires that the wavelength-dependent, spectral directive gain density be defined over the wavelength range of the normalized radiation intensity spectral density,
Note that is normalized over the total spectrum, where the largest value of is not equal to 1, which was the case for . The normalization condition results in after integration of over the wavelength spectrum. The average normalized radiation intensity spectral density integrated over the sphere is given by
It is often difficult for experimental data to capture the full range of a theoretical concept. For example, no single detector can be used for all possible frequencies ranging from radio frequency to x-rays. Therefore, measurement data must account for these experimental limitations. For externally pumped samples, the detector cannot be allowed to block the pump beam. Therefore, low co-latitude angles are not always available to scan. Likewise, the apparatus presented in this article has a sample cooling stage, which is useful for biological materials that can degrade at a higher rate at room temperature. Therefore, the sample mounting technique will also affect the range over which the emission can be scanned. The average normalized radiation intensity spectral density integrated over the sphere given in Equation 11 cannot be divided by the solid angle of a full sphere if the scanning apparatus could not scan for light. In some cases, the intensity can be assumed to be quite low in some directions such as when a sample is mounted on a mirrored surface. Other cases may indeed allow for much light to be emitted, but unable to detect with the apparatus. Therefore, we must consider a worthy measure over the scanned surface.
Let us consider the integral in the denominator of Equation 11. If the scanned area is not over the full sphere, then we should consider normalizing the function over the scanned area only. Then substituting the normalized radiation intensity spectral density integrated over the scanned area only, we arrive at a metric useful for incomplete scans, which is the scanned-area, spectral directive gain density,
where the denominator is averaged over integral limits associated with the angular patch of the sphere that was scanned by the apparatus. Integrating over all wavelengths yields the scanned-area, directive gain, .
Note that the total power emitted over the entire spectrum can be useful to find hot spots in light sources that are relatively unremarkable in their spectrum as a function of angle. Examples of such light sources are light-emitting diodes (LED), incandescent light bulbs, and thin films doped with a low-concentration of fluorophores. The apparatus presented in this article collects raw spectral data, which can be used to determine the spectral directive gain. The alignment is characterized by generating visual map of the directive gain while the wavelength-dependent, spectral directive gain is visualized at specific angles to study the dominance of different emission processes directed at different angles.
III Experimental methods
III.1 Apparatus
@@
Automated recordings of emission spectra scanned over the azimuth and polar angles were performed by an in-house built apparatus. An aluminum sample stage was at the center of the apparatus and temperature controlled by running liquid through a bored cavity. The sample stage was secured by a post that ran through a bore hole in a rotation stage. The sample stage remained in a fixed position while the rotation stage was allowed to sweep the azimuth angle as shown in Figure 1. A circular plate was then attached to the azimuth rotation stage. An off-center rotation stage was attached to the circular plate so that it could sweep the polar angles. The circular plate was used to mark rotation angles to initially configure the azimuth rotation stage at the speed of /s. A boom arm was attached to the second rotation stage which swept the polar angle. The polar angle was displaced between two positions at a speed of /s. A lever arm was attached at the top of the boom arm to secure the detector and its alignment components.
@@
A tunable double lens tube was used to zoom and focus in on the sample stage. Adjustments to the lens tube’s orientation were made by mounting it on a kinematic mirror mount. The kinematic mirror mount is mounted on a slot and the distance between zoom lens tube and the sample plate can be adjusted, A camera was placed in the image location for alignment purposes. After alignment, the camera was replaced by a m-core multimode optical patch fiber. The fiber was connected to a portable spectrometer mounted to the boom arm. We initially attempted to secure one end of a long fiber to the apparatus and the other end to a spectrometer mounted on the optical table, but fiber deformations from apparatus rotations caused the relative intensity of the signal to change. By mounting the spectrometer to the boom arm with a rigid fiber patch cord, loss through the optical fiber was held constant. A slot was located between the lens tube and the detector position which allowed for 1” filters (chromatic, polarization, etc.).
@@
The motors on the apparatus were controlled by a CNC shield attached to an Arduino Uno loaded with a GRBL library. The Arduino was connected to a computer via USB port. The spectrometer was also connected to the computer. A Python script recorded spectra with the Seabreeze package, and it also controlled the apparatus motors. A third motor was connected to the CNC shield which rotated a half-wave plate as shown in Figure 1. The pump beam’s polarization was rotated and sent through a polarizing beam cube splitter. Pulse energy was adjusted using this half-wave plate and polarizer combination. A relay was also connected by USB port and incorporated into the same Python code that controlled the entire apparatus. The relay controlled a shutter, where the shutter was open for data collection of materials excited by the external pump laser and closed when the apparatus rotated during the scanning procedure. The shutter reduced fluorophore photodegradation during a scan.
III.2 Alignment
@@ III.3 Apparatus testing and experimentation
The apparatus measures the directive gain of an emissive material/device over the spectral range recorded by the fiber spectrometer. Two types of samples were recorded, 1) an electrically-driven light emitting diode (LED) with a diffuser cap and 2) a fluorescein-doped water droplet. The electrical current for the LED and the pump beam pulse energy for the fluorescein-doped water were held constant for the duration of their scans. The droplets were carefully measured before each scan and after each scan using a plastic scale that was placed underneath the droplet. All scans were stepped in the azimuth from to . They scans were stepped along the co-latitude from to .
@@
The two different sizes of droplet samples were measured under the same geometric condition. For the first one, a large dye-doped water droplet was placed on a parafilm surface with a pump beam illuminating the sample from above. The hydrophobic surface caused the droplet to form a dome shape. A second droplet’s diameter was approximately half the diameter of the first droplet and pumped with a similar size pump beam. A nm-wavelength, ns pulsed laser beam operating at Hz was used to pump the liquid droplets. The wavelength was chosen as nm because it is highly absorbed by fluorescein while also absorbed by GG475 Schott glass. The GG475 filter placed in the detector optics chain passed all detectable fluorescence while blocking stray pump light that can saturate spectrometer pixels.
IV Results
IV.1 Diffuser-capped LED
@@
Our system was aligned through the process described in Section III.2. In an attempt to confirm the alignment, an LED with a diffuser cap was placed at the center of the sample stage. An ideal spherical diffuser cap would mimic a spherically symmetric light source, which would confirm the alignment; however, the spherical diffuser capping the LED did not have spherical light-emissive symmetry. Although the normalized spectral profile was independent of angle, the output intensity had angular dependencies along both the polar and azimuth directions.
@@
At high polar angles, a hot spot was formed along the azimuth as depicted in the heat map representing the scanned-area, spectral directive gain density given in Figure 2(a). The nonuniform could have been caused by an alignment issue, so the sample was rotated by and a new scan was performed. As shown in Figure 2(b), the recorded hot spot rotated in the azimuthal direction by the same angle that the sample was rotated. This azimuth rotation of the sample confirmed the hot spot was indeed sourced from the diffuser-capped LED as opposed to an issue with the system alignment.
@@
As expected, there was no distinguishable difference between the peak wavelength and spectral width of the LED’s emitted light. The amplitude of the spectra shown in Figure 2(c-f) vary in amplitude only. There is a significant trend for the dependence on the polar angle as shown in Figure 2(c-d). The detected signal is largest near the top of the LED and significantly weakens as the detector collects light near the horizon. A hot spot is clearly identifiable as a function of the azimuthal angle, the peak amplitude of the spectra, represented by and shown in Figure 2(e-f), shifted by the same angle as the rotated sample. The LED has no significant process for emitting amplified light, and the diffuser cap is fabricated from a passive achromatic material. Therefore, as expected for a linear optical source, the spectral features remain constant while only the amplitude of the emitted light depends on the spherical angles. This result implies that highly directed energy is emitted from the source or during the passive diffusion through the diffuser cap, as expected.
IV.2 Large water droplet: center pumped
@@
The diffuser-capped, electrically-pumped, LED light source in the previous section had hot spots caused by aspherical light emission. A wt.% fluorescein/water solution was also prepared in an attempt to create a better approximation to a spherical-emitting light source. The droplet was placed on parafilm, where the solution/air interface above the parafilm surface formed an dome that can be approximated as a hemisphere. The beam spot was in diameter, which illuminated the majority of the liquid droplet’s surface. The liquid cooled stage was measured to be C while the ambient air was held at room temperature and moderate humidity. The cooled surface resulted in a near steady state droplet size due to the competition between evaporation and condensation.
@@
There is an increased emission intensity at higher co-latitude (lower latitude) for the droplet. This trend in emission as a function of polar angle is opposite of the diffuser-capped LED, where the scanned-area, spectral directive gain density shown in Figure 3(a) can be visually compared to those given in Figure 2(c).
@@
The noise associated with the spectrum is also greater in the droplet at low emission intensity levels as compared to the diffuser-capped LED. The azimuthal dependence shown in Figure 2(b) also has some shot-to-shot variations throughout the scan. Some small variations in the beam profile could result in an azimuth-dependent emission; however,noise from the pump beam also contributes to the variations in the azimuthal dependence of the emission intensity. The shot-to-shot noise of an Nd:YAG’s fundamental frequency is reasonable low. The shot-to-shot uncertainty of the pump is increased when a portion of the nm pump beam’s photons are frequency doubled by a beta Barium Borate (BBO) crystal followed by sum-frequency generation of the nm and nm to yield a nm beam. This nonlinear, macroscopic, cascading process results in larger variations between pulse energies for the nm beam entering an optical parametric oscillator (OPO). Prior to the pump light entering the OPO cavity, the residual nm and nm light is attenuated by a set of dichroic mirrors. The remaining nm light is injected into a cavity via another dichroic mirror with a long BBO crystal between a high reflectance mirror and a dichroic output coupler. The amplified output of the OPO for wavelength selection at nm also amplifies the shot-to-shot uncertainty of the pulse energy. Thus, the uncertainty between each shot appears to be relatively large.
@@
Averaging out the uncertainty between recorded shots taken while the azimuth is scanned, a slow change of intensity is observed through the scan as shown in Figure 3(b). Therefore, the calculated shows some preference for an azimuthal emission direction. If the preferred direction were biaxial, the cause would undoubtedly be from the pump beam’s polarization. The excited state lifetime of fluorescein is relatively long, making a polarization dependent emission unlikely in water. Indeed, we found that the small, preferred, azimuth emission was unidirectional, which was caused by a small uncertainty in both the pump beam profile and center-alignment. The ASE clearly dominates the fluorescence signal. The small fluorescence peak is observed at a nm wavelength while the peak of the red-shifted ASE profile is much greater at nm.
IV.3 Small water droplet: off-center pumped